Given the two non-linear equations: fi(x, y) = y - x and f2(x, y) = 36 – 4x2 +9y2 With an initial approximation of (1,2), finHand calculations please !
Answer #1

Solution:
f1=y-c=0→(1)

f2=9y2-4b+36=0→(2)

Here total variable is 3(b,c,y), but total equation is 2, so solution is not possible